The Heaths are paying off two car loans. One loan charges 5% interest per year. The other loan charges 8% interest per year. They owe $2400 more on the 5% interest loan than they do on the other. Last year they paid a total of $710.20 interest. How much do they owe on each loan?
Answer by mananth(16946)

rate------amount----interest
5% ------- x -------0.05x
8%---------y--------0.08y
x=y+2400
x-y=2400............1
0.05x+0.08y=710.20...2
x-y=2400 .............1
0.05x+0.08y=710.2.............2
Eliminate y
multiply (1)by 0.08
Multiply (2) by 1
0.08x-0.08y=192
0.05x+0.08y=710.20
Add the two equations
0.13 x = 902.2
/0.13
x= 6,940.00
plug value of x in (1)
x-y=2400
6940-y=2400
-y=2400-6940
-y=-4540
y=4540.00
m.ananth@hotmail.ca